OpenStudy (anonymous):

please help. 2 csc x-3=-5 on the interval (0,2pi)

OpenStudy (e.mccormick):

What have you tried on this? Perhaps a little algebra to get \(\csc x\) alone?

OpenStudy (anonymous):

$$2\csc x-3=-5\\2\csc x=-5+3=-2\\\csc x=-1\\\frac1{\sin x}=-1\\\sin x=-1\\x=?$$

OpenStudy (anonymous):

-1. i understand how to get that far but how do i take that answer and put it on the unit circle

OpenStudy (e.mccormick):

Well, where is the sine value -1?

OpenStudy (e.mccormick):